# DEFOM-Stereo: Depth Foundation Model Based Stereo Matching

### v1

@@ Abstract
Stereo matching is a key technique for metric depth estimation in computer vision and robotics. Real-world challenges like occlusion and non-texture hinder accurate disparity estimation from binocular matching cues. Recently, monocular relative depth estimation has shown remarkable generalization using vision foundation models. Thus, to facilitate robust stereo matching with monocular depth cues, we incorporate a robust monocular relative depth model into the recurrent stereo-matching framework, building a new framework for depth foundation model-based stereo-matching, DEFOM-Stereo. In the feature extraction stage, we construct the combined context and matching feature encoder by integrating features from conventional CNNs and DEFOM. In the update stage, we use the depth predicted by DEFOM to initialize the recurrent disparity and introduce a scale update module to refine the disparity at the correct scale. DEFOM-Stereo is verified to have comparable performance on the Scene Flow dataset with state-of-the-art (SOTA) methods and notably shows much stronger zero-shot generalization. Moreover, DEFOM-Stereo achieves SOTA performance on the KITTI 2012, KITTI 2015, Middlebury, and ETH3D benchmarks, ranking 1st on many metrics. In the joint evaluation under the robust vision challenge, our model simultaneously outperforms previous models on the individual benchmarks. Both results demonstrate the outstanding capabilities of the proposed model.
@@ 1 Introduction
Stereo matching involves recovering the disparity between two views and has been a fundamental topic in computer vision for decades. Stereo matching has obtained steady improvement with the consistently proposed algorithms, which range from hand-designed ones [38] to end-to-end deep architectures [22, 46]. The deep matching frameworks have mainly shifted from cost-volume filtering ones [21, 5, 15, 56, 64] to recent recurrent refinement ones [27, 23, 67, 54, 7, 50]. According to [46], despite RAFT-Stereo [27] has made good progress, the generalization and robustness remain challenging. The critical challenges lie in occlusion, non-texture, image blur, high-resolution, etc.
@@
On the other hand, a closely related task, monocular relative depth estimation [34, 35, 2, 20, 60, 61] has achieved advancing zero-shot generalization ability across diverse scenes. Among them, Depth Anything V2 [61], is a model with a pre-trained vision transformer (ViT) [9] backbone of the general-purpose vision foundation model DINOv2 [33] and a head of the dense prediction transformer (DPT) [35]. It is trained on synthetic labeled images and distilled on large realistic unlabeled images to conduct robust and efficient depth prediction and recover impressive fine-grained details. These properties allow it to be a depth foundation model (DEFOM). Nevertheless, DEFOM does not provide metric depth, which is crucial for robotic applications. To achieve more robust stereo depth estimation, we incorporate Depth Anything V2 into RAFT-Stereo [27] to build a DEFOM-based stereo matching (DEFOM-Stereo) model.
| Sceneflow | CREStereo | KITTI-2015 | Middlebury-half | ETH3D | |||||
| EPE | STD | EPE | STD | EPE | STD | EPE | STD | EPE | STD |
| 8.04 | 3.31 | 5.10 | 3.30 | 2.08 | 0.74 | 5.00 | 0.11 | 0.65 | 0.40 |
@@
We first consider using DEFOM to improve the feature extraction for stereo matching. Utilizing the pre-trained ViT to improve stereo matching has recently been explored in [29, 66]. Both focus on developing attention-based adapters to convert the ViT backbone features into matching features. In contrast, we use a DEFOM that provides dense feature maps with a DPT head and demonstrates that a simple convolutional fusion of DPT features and plain CNN features is effective. As the existing DPT should be fixed to predict depth, we initialize a trainable DPT head to provide more flexible features for fusion. Apart from the matching feature, there is a context encoder of the recurrent stereo framework to provide monocular cues controlling the recurrent disparity update process. Thus, we propose a combined matching feature and context encoder with CNN and DEFOM.
@@
Apart from its features, the DEFOM depth can intuitively be used to improve disparity recovery. The DEFOM depth is an affine disparity with unknown scale and shift. We first assess the depth with the metric disparity of typical stereo datasets using the least-square affine alignment. We compute the end-point error (EPE) between the aligned depth and ground-truth disparity, and the standard deviation (STD) of their ratios (details in supplement) in Tab. 1. Even given the unknown scale and shift, disparity errors are large, especially for the synthetic Scene Flow [30] and CREStereo [23] datasets, whose STD is very high too, indicating the scale inconsistency within the image is serious. This is because the synthetic stereo dataset is about unnatural scenes, such as an example in Scene Flow shown in Fig. 4. As the stereo model is usually pre-trained on the synthetic stereo dataset, the scale inconsistency of DEFOM poses challenges for recovering disparity from its depth estimate.
@@
To cope with the scale inconsistency, we propose the scale update (SU) module that performs recurrent dense scale updates on the initialized disparity from the depth estimate. We found that the depth amplitude by Depth Anything V2 varies across different image samples with different backbone sizes. Considering that the disparity range is usually proportional to image resolution, we use the image width to normalize the depth estimate in initialization. To facilitate SU, we design a scale lookup (SL) on the correlation volume on which the current disparity estimate is multiplied with a set of scale values to generate sampling indexes. Compared with the conventional pyramid lookup (PL), SL has a complete search range on the whole image, enabling SU with global matching. Therefore, we insert the SU module before the conventional delta update (DU) module in RAFT-Stereo [27], which can continue to recover local details upon the result of SU.
@@
The overall framework of DEFOM-Stereo is illustrated in Fig. 5. We have conducted extensive experiments to verify DEFOM-Stereo’s effectiveness. When pre-training on Scene Flow, DEFOM-Stereo achieves comparable performance with SOTA methods and obtains considerable progress in the zero-shot generation evaluation on KITTI 2012 and 2015, Middlebury, and ETH3D, for example, reducing the 2-pixel thresholding error rate by over a half on Middlebury. When submitting to the official benchmarks of these datasets, DEFOM-Stereo achieves top performance, ranking on many metrics on the four leaderboards among all submissions while writing this paper. A brief comparison with published SOTA methods is demonstrated in Fig. 2, where DEFOM-Stereo exhibits remarkable advantages. Furthermore, we also examine our model under the joint evaluation of the Robust Vision Challenge (RVC), and our model simultaneously outperforms all previous RVC models on the three benchmarks, particularly on KITTI 2015 and Middlebury with a large margin.
@@
Our model achieves large performance gains on synthetic-to-real zero-shot generalization compared to SOTA models and is top-ranking on typical stereo leaderboards of KITTI, Middlebury, and ETH3D simultaneously.
@@ 3.7 Loss Function
Suppose that the total number of update iterations is and the set of predicted disparity maps are . We follow [45, 27] to supervise those predictions in exponentially increasing weights with the ground truth disparity . Tho loss is computed as,
| Methods | RAFT-Stereo [27] | DLNR [67] | IGEV-Stereo [54] | NMRF-Stereo [14] | Selective-IGEV [50] | Mocha-Stereo [7] | DEFOM-Stereo (Ours) |
| EPE [px] | 0.56 | 0.48 | 0.47 | 0.45 | 0.44 | 0.41 | 0.42 |
| Bad 1.0 [%] | 6.63 | 5.39 | 5.21 | 4.50 | 4.98 | 5.07 | 5.10 |
@@ 4 Experiments
The experiments performed common datasets, including the synthetic Scene Flow [30], and realistic KITTI [13, 31], Middlebury [39] and ETH3D [40]. Initially, we trained the model on Scene Flow and compared it with SOTA methods on in-domain fitting and cross-domain generalization. Next, the effectiveness of the proposed components is verified through an ablation study. Finally, a comparison of the official benchmarks of the realistic datasets is presented.
4.1 Implementation Details
We implement our DEFOM-Stereo using the PyTorch framework on NVIDIA RTX 4090 GPUs. We use a simplified REAF-Stereo that reduces the correlation pyramid level from 4 to 2 as our baseline model and add the proposed components to construct DEFOM-Stereo. For all training, we use the AdamW optimizer, a one-cycle learning rate schedule with a learning rate of 2e-4 and a batch size of 8. The total number of update iterations (SU+DU) is set to 18 in training and 32 in evaluation. The number of SU is fixed to 8 during both training and evaluation.
4.2 Preliminary Comparisons
In this section, following previous work [27, 67, 54, 14, 50, 7], we first pre-train our model on Scene Flow for 200k. In pre-training, the crop size is set to .
Pre-training Evaluation. We first make a comparison with SOTA methods on the test set of Scene Flow to examine the fitting ability of our model. The results are listed on Tab. 2. DEFOM-Stereo has achieved comparable performance with recent SOTA methods [14, 50, 7] and reduces EPE by compared with the RAFT-Stereo.
Zero-shot Generalization. We follow [64, 27] to examine the zero-shot generalization ability of our model on the trainsets of four realistic datasets, and for Middlebury, we consider the three available resolutions. For Middlebury and ETH3D, we count all valid pixels instead of only those in non-occluded regions, as this allows for a more comprehensive test. For other methods, we perform such evaluation using the provided pre-trained model on Scene Flow. The results are shown in Tab. 3. We found that despite later recurrent optimation-based methods [67, 54, 50, 7] have improved on pre-training fitting on Scene Flow compared with RAFT-Stereo, fail to make comprehensive progress on zero-shot evaluation. For example, DLNR [67], shows a clear performance gain on Middlebury with different resolutions, but significantly increases the error rate on KITTI and ETH3D. For the recent neural Markov random field-based model, NMRF-Stereo [14], progress can be seen only in driving datasets KITTI 2012 and 2015, and Middlebury-quarter.
@@ 4.2 Preliminary Comparisons
| Mocha-Stereo [7] | 4.85 | 5.93 | 33.42 | 9.11 | 4.66 | 3.89 |
| DEFOM-Stereo (ViT-S) | 4.29 | 5.29 | 10.76 | 4.29 | 3.67 | 2.61 |
| DEFOM-Stereo (ViT-L) | 3.76 | 4.99 | 8.51 | 3.26 | 3.27 | 2.35 |
4.3 Ablation Study
@@
| +CCE | ✓ | 0.49 | 6.08 | 4.40 | 5.84 | 5.67 | 2.82 | 12.10 | 0.242 | |||
| +CFE | ✓ | 0.50 | 6.17 | 4.13 | 5.53 | 7.55 | 2.83 | 13.89 | 0.243 | |||
| +CCE+CFE | ✓ | ✓ | 0.49 | 5.95 | 4.02 | 5.75 | 5.56 | 2.53 | 14.11 | 0.246 | ||
| +DI | ✓ | 0.57 | 6.74 | 4.57 | 5.63 | 9.40 | 2.77 | 11.11 | 0.242 | |||
| +DI+SU | ✓ | ✓ | 0.50 | 6.03 | 4.15 | 5.12 | 5.23 | 2.67 | 15.51 | 0.244 | ||
| Full Model (ViT-S) | ✓ | ✓ | ✓ | ✓ | 0.46 | 5.57 | 4.29 | 5.29 | 4.29 | 2.61 | 18.51 | 0.255 |
| Full Model (ViT-L) | ✓ | ✓ | ✓ | ✓ | 0.42 | 5.10 | 3.76 | 4.99 | 3.26 | 2.35 | 47.30 | 0.316 |
In this section, we verify the effectiveness of the proposed components via a main ablation study. The supplementary material will present a detailed ablation study of the combined encoders’ design choices and the proposed scale update module’s number of steps. In the main ablation study, we mainly use the ViT-S as the ViT backbone for our model and train all the variants on Scene Flow for 200k steps. For accuracy comparison, both the in-domain test and zero-shot generalization evaluation are presented. We also list trainable parameters and the inference time of the model variants for comparing computational complexities. The results are shown in Tab. 4.
@@ Appendix A Evaluation of Depth Anything V2
In this section, we present the details of the evaluation of Depth Anything V2 [61] on some common stereo datasets. To achieve cross-scene robustness, the relative depth estimation models are usually trained with a scale and shift (affine) invariant loss [34] on the inverse depth space, thus predicting an affine disparity with unknown scale and shift. Given an inverse depth map predicted by Depth Anything V2 is and its corresponding ground truth disparity map , they must satisfy the following affine transformation,
@@
In the introduction section of the main paper, we have presented the results of three realistic datasets, KITTI 2015 [31], Middlebury (half resolution) [39] and ETH3D [40], and two synthetic datasets, Scene Flow [30] and CREStereo [23]. For the realistic datasets, we evaluate on their entire trainsets, i.e., 200 examples for KITTI 2015, 15 samples for Middlebury, and 27 samples for ETH3D. For Scene Flow and CREStereo, we evaluate on 200 random samples from their trainsets. The results are shown in Table 1 of the main paper. Depth Anything V2 presents good results on realistic datasets. Although the EPE on Middlebury-half is up to 5, it is mainly due to its high resolution. In contrast, for two synthetic datasets, both the EPE and STD are larger, indicating Depth Anything V2 does not predict depth maps with good scale consistency.
@@ B.2 Iterations of Scale Update
In this section, we investigate the effect of the number of iterations of the scale update module. Likewise, we fixed the total number of iterations as 18 in training and 32 in evaluation, and the number of SU iterations is set the same in training and evaluation. We perform training on Scene Flow for 50k steps with a batch size of 4 and evaluation on the Scene Flow test set. The number of scale update iterations is increased from 0 to 10 and the 0 scale update iteration represents that the scale update module is not used. When the scale update iteration is 0, the model has the highest error metrics, and increasing it to 1 results in over error reduction. The EPE continues to decrease until the scale update iteration reaches 8 and the performance for the 7-9 scale update iteration is closed. When the scale update iteration exceeds 9, the performance starts to drop obviously. Therefore, we select 8 as the number of scale update iterations in our final model.

### v2

@@ Abstract
Stereo matching is a key technique for metric depth estimation in computer vision and robotics. Real-world challenges like occlusion and non-texture hinder accurate disparity estimation from binocular matching cues. Recently, monocular relative depth estimation has shown remarkable generalization using vision foundation models. Thus, to facilitate robust stereo matching with monocular depth cues, we incorporate a robust monocular relative depth model into the recurrent stereo-matching framework, building a new framework for depth foundation model-based stereo-matching, DEFOM-Stereo. In the feature extraction stage, we construct the combined context and matching feature encoder by integrating features from conventional CNNs and DEFOM. In the update stage, we use the depth predicted by DEFOM to initialize the recurrent disparity and introduce a scale update module to refine the disparity at the correct scale. DEFOM-Stereo is verified to have much stronger zero-shot generalization compared with SOTA methods. Moreover, DEFOM-Stereo achieves top performance on the KITTI 2012, KITTI 2015, Middlebury, and ETH3D benchmarks, ranking on many metrics. In the joint evaluation under the robust vision challenge, our model simultaneously outperforms previous models on the individual benchmarks, further demonstrating its outstanding capabilities.
@@ 1 Introduction
Stereo matching involves recovering the disparity between two views and has been a fundamental topic in computer vision for decades. Stereo matching has obtained steady improvement with the consistently proposed algorithms, which range from hand-designed ones [40] to end-to-end deep architectures [24, 48]. The deep matching frameworks have mainly shifted from cost-volume filtering ones [23, 6, 17, 59, 67] to recent recurrent refinement ones [29, 25, 70, 57, 8, 52]. According to [48], despite RAFT-Stereo [29] has made good progress, the generalization and robustness remain challenging. The critical challenges lie in occlusion, non-texture, image blur, high-resolution, etc. etc.
@@
On the other hand, a closely related task, monocular relative depth estimation [36, 37, 3, 22, 63, 64] has achieved advancing zero-shot generalization ability across diverse scenes. Among them, Depth Anything V2 [64], is a model with a pre-trained vision transformer (ViT) [11] backbone of the general-purpose vision foundation model DINOv2 [35] and a head of the dense prediction transformer (DPT) [37]. It is trained on synthetic labeled images and distilled on large realistic unlabeled images to conduct robust and efficient depth prediction and recover impressive fine-grained details. These properties allow it to be a depth foundation model (DEFOM). Nevertheless, DEFOM does not provide metric depth, which is crucial for robotic applications. To achieve more robust stereo depth estimation, we incorporate Depth Anything V2 into RAFT-Stereo [29] to build a DEFOM-based stereo matching (DEFOM-Stereo) model. To be specific, we perform the incorporation with two aspects, 1) utilizing DEFOM’s pre-trained features to improve the feature extraction for stereo matching, and 2) facilitating disparity updating with the predicted monocular depth.
@@
Utilizing the pre-trained ViT to improve stereo matching has recently been explored in [31, 69] via developing attention-based adapters to convert the ViT backbone features into matching features. In contrast, the DEFOM provides dense feature maps with a DPT head, and the feature is more relevant to the depth task. We demonstrate that a simple convolutional fusion of DPT features and plain CNN features is effective. As the existing DPT should be fixed to predict depth, we initialize a trainable DPT head to provide more flexible features for fusion. Apart from the matching feature, there is a context encoder of the recurrent stereo framework to provide monocular cues controlling the recurrent disparity update process. Thus, we propose a combined matching feature and context encoder with CNN and DEFOM.
@@
Besides its features, the DEFOM depth can intuitively be used to improve disparity recovery. The DEFOM depth is an affine disparity with unknown scale and shift. However, finding a single scale and shift values does help to recover accurate disparity. A preliminary evaluation of Depth Anything V2 [64] on typical stereo datasets using the least-square affine alignment illustrates the disparity error is significantly large, indicating the scale inconsistency for depth prediction within the image is serious. The phenomenon is more serious for synthetic stereo datasets, for example of the synthetic Scene Flow [32] shown in Fig. 3. We conjecture that the scale inconsistency is caused by the training in different FoV images with affine invariant loss. As the stereo model is usually pre-trained on the synthetic stereo dataset, the scale inconsistency of DEFOM poses challenges for recovering disparity from its depth estimate.
@@
To cope with the scale inconsistency, we propose a scale update (SU) module that performs recurrent dense scale updates on the initialized disparity from the depth estimate. We found that the depth amplitude by Depth Anything V2 varies across different image samples with different backbone sizes. Considering that the disparity range is usually proportional to image resolution, we use the image width to normalize the depth estimate in initialization. To facilitate SU, we design a scale lookup (SL) on the correlation volume on which the current disparity estimate is multiplied with a set of scale values to generate sampling indexes. Compared with the conventional pyramid lookup (PL), SL has a complete search range on the whole image, enabling SU with global matching. Therefore, we insert the SU module before the conventional delta update (DU) module in RAFT-Stereo [29], which can continue to recover local details upon the result of SU.
@@
The overall framework of DEFOM-Stereo is illustrated in Fig. 4. We have conducted extensive experiments to verify DEFOM-Stereo’s effectiveness. When pre-training only on Scene Flow, DEFOM-Stereo achieves comparable performance with SOTA methods and obtains considerable progress in the zero-shot generation evaluation on KITTI 2012 and 2015, Middlebury, and ETH3D, for example, reducing the 2-pixel thresholding error rate by about a third on Middlebury. When submitting to the official benchmarks of these datasets, DEFOM-Stereo achieves top performance, ranking on many metrics on the four leaderboards among all submissions while writing this paper. A brief comparison with published SOTA methods is demonstrated in Fig. 2, where DEFOM-Stereo exhibits remarkable advantages. Furthermore, we also examine our model under the joint evaluation of the Robust Vision Challenge (RVC), and our model simultaneously outperforms all previous RVC models on the three benchmarks, particularly on KITTI 2015 and Middlebury with a large margin.
@@ 2.2 Zero-shot Generalized Stereo Matching
Note that there are some contemporary works [2, 9, 55] that incorporate the depth foundation model into the recurrent stereo matching framework. They both use 3D convolutions upon a 4D cost volume to obtain initial disparity following IGEV-Stereo [57] and [2, 9] further get another refined initial disparity via alignment with the DEFOM depth. Our model differs in directly replacing the initial disparity of the original RAFT-stereo with the DEFOM depth and reforming the recurrent update process accordingly.
@@ 4 Experiments
In this section, we verify the effectiveness of our model via synthetic-to-real zero-shot generalization, ablation study and evaluation on popular online stereo benchmarks. The main ablation study shows the effectiveness and complexities of the proposed components and the additional ablation study of the combined encoders’ design choices and the iterations of scale updates will be presented on supplementary.
@@
| Mocha-Stereo [8] | 4.85 | 5.93 | 35.34 | 11.49 | 7.39 | 3.89 |
| DEFOM-Stereo (ViT-S) | 4.29 | 5.29 | 14.70 | 6.76 | 6.38 | 2.61 |
| DEFOM-Stereo (ViT-L) | 3.76 | 4.99 | 11.95 | 5.91 | 5.65 | 2.35 |
4.1 Implementation Details
We implement our DEFOM-Stereo using the PyTorch framework on NVIDIA RTX 4090 GPUs. We use a simplified REAF-Stereo that reduces the correlation pyramid level from 4 to 2 as our baseline model and add the proposed components to construct DEFOM-Stereo. For all training, we use the AdamW optimizer, a one-cycle learning rate schedule with a learning rate of 2e-4 and a batch size of 8. The total number of update iterations (SU+DU) is set to 18 in training and 32 in evaluation. The number of SU is fixed to 8 during both training and evaluation. in Eqn. 4 is set to 0.05.
4.2 Zero-shot Generalization
We first pre-train our model on Scene Flow for 200k with a crop size of . We follow [67, 29] to examine the zero-shot generalization ability of our model on the trainsets of four realistic datasets, and for Middlebury, we consider the three available resolutions. For Middlebury and ETH3D, we count all valid pixels instead of only those in non-occluded regions, as this allows for a more comprehensive test. For other methods, we perform such evaluation using the provided pre-trained model on Scene Flow. The results are shown in Tab. 1. We found that despite later recurrent optimation-based methods [70, 57, 52, 8] have improved on pre-training fitting on Scene Flow compared with RAFT-Stereo, fail to make comprehensive progress on zero-shot evaluation. For example, DLNR [70], shows a clear performance gain on Middlebury with different resolutions, but significantly increases the error rate on KITTI and ETH3D. For the recent neural Markov random field-based model, NMRF-Stereo [16], progress can be seen only in driving datasets KITTI 2012 and 2015, and Middlebury-quarter.
@@ 4.2 Zero-shot Generalization
| +CCE | ✓ | 0.49 | 6.08 | 4.40 | 5.84 | 8.42 | 2.82 | 12.10 | 0.242 | |||
| +CFE | ✓ | 0.50 | 6.17 | 4.13 | 5.53 | 10.45 | 2.83 | 13.89 | 0.243 | |||
| +CCE+CFE | ✓ | ✓ | 0.49 | 5.95 | 4.02 | 5.75 | 8.31 | 2.53 | 14.11 | 0.246 | ||
| +DI | ✓ | 0.57 | 6.74 | 4.57 | 5.63 | 12.40 | 2.77 | 11.11 | 0.242 | |||
| +DI+SU | ✓ | ✓ | 0.50 | 6.03 | 4.15 | 5.12 | 8.15 | 2.67 | 15.51 | 0.244 | ||
| Full Model (ViT-S) | ✓ | ✓ | ✓ | ✓ | 0.46 | 5.57 | 4.29 | 5.29 | 6.76 | 2.61 | 18.51 | 0.255 |
| Full Model (ViT-L) | ✓ | ✓ | ✓ | ✓ | 0.42 | 5.10 | 3.76 | 4.99 | 5.91 | 2.35 | 47.30 | 0.316 |
4.3 Ablation Study
In this section, we verify the effectiveness of the proposed components via a main ablation study. We mainly use the ViT-S as the ViT backbone for our model and train all the variants on Scene Flow for 200k steps. For accuracy comparison, both the in-domain test and zero-shot generalization evaluation are presented. We also list trainable parameters and the inference time of the model variants for comparing computational complexities. The results are shown in Tab. 2.
@@ Appendix A Evaluation of Depth Anything V2
In this section, we present the evaluation of Depth Anything V2 [64] on some common stereo datasets. To achieve cross-scene robustness, the relative depth estimation models are usually trained with a scale and shift (affine) invariant loss [36] on the inverse depth space, thus predicting an affine disparity with unknown scale and shift. Given an inverse depth map predicted by Depth Anything V2 is and its corresponding ground truth disparity map , they must satisfy the following affine transformation,
@@
We perform evaluation on three realistic datasets, KITTI 2015 [33], Middlebury (half resolution) [41] and ETH3D [42], and two synthetic datasets, Scene Flow [32] and CREStereo [25]. For the realistic datasets, we evaluate on their entire trainsets, i.e., 200 examples for KITTI 2015, 15 samples for Middlebury, and 27 samples for ETH3D. For Scene Flow and CREStereo, we evaluate on 200 random samples from their trainsets.
Tab. 6 presents the quantitative results. Even given the unknown scale and shift, disparity errors are large, especially for the synthetic Scene Flow [32] and CREStereo [25] datasets, whose STD is very high too, indicating the scale inconsistency within the image is serious. This is because the synthetic stereo dataset is about unnatural scenes. As the stereo model is usually pre-trained on the synthetic stereo dataset, the scale inconsistency of DEFOM poses challenges for recovering disparity from its depth estimate. In contrast, Depth Anything V2 presents better results on realistic datasets. Although the EPE on Middlebury-half is up to 5, it is mainly due to its high resolution. In contrast, for two synthetic datasets, both the EPE and STD are larger, indicating Depth Anything V2 does not predict depth maps with good scale consistency.
@@ B.2 Iterations of Scale Update
In this section, we investigate the effect of the number of iterations of the scale update module in Tab. 8. Likewise, we fixed the total number of iterations as 18 in training and 32 in evaluation, and the number of SU iterations is set the same in training and evaluation. We perform training on Scene Flow for 50k steps with a batch size of 4 and evaluation on the Scene Flow test set. The number of scale update iterations is increased from 0 to 10 and the 0 scale update iteration represents that the scale update module is not used. When the scale update iteration is 0, the model has the highest error metrics, and increasing it to 1 results in over error reduction. The EPE continues to decrease until the scale update iteration reaches 8 and the performance for the 7-9 scale update iteration is closed. When the scale update iteration exceeds 9, the performance starts to drop obviously. Therefore, we select 8 as the number of scale update iterations in our final model.